OpenStudy (anonymous):

How to Define the following functions? help please.. 1. f(x)=x-5; g(x)= x^2-1 a. f+g b. f-g c. f⋅g d. f/g e. g/f

OpenStudy (anonymous):

a. f+g (x-5)+(x^2-1) = x^2+x-6 b. f-g (x-5)-(x^2-1) = -x^2+x-4

OpenStudy (anonymous):

@mathslover help here please..

OpenStudy (ranga):

You have got the first two right. Just continue on. Multiply the two functions for c. Divide f by g for d and take its reciprocal for e

OpenStudy (anonymous):

@ranga how to solve for multiplication and division ??

OpenStudy (ranga):

f.g = (x-5)(x^2 - 1) Expand the above.

OpenStudy (anonymous):

can you show me how to expand it??

OpenStudy (anonymous):

how if we use the FOIL method in the multiplication ??

OpenStudy (ranga):

Use FOIL.

OpenStudy (ranga):

(a+b)(c+d) = ac + ad + bc + bd

OpenStudy (ranga):

(x-5)(x^2 - 1) = x^3 -x -5x^2 + 5 = x^3 - 5x^2 - x + 5

OpenStudy (ranga):

f/g = (x - 5)/(x^2 - 1)

OpenStudy (anonymous):

okok

OpenStudy (anonymous):

i dont know the proess if it is in division..

OpenStudy (anonymous):

@ranga still there ??

OpenStudy (ranga):

f/g = (x - 5)/(x^2 - 1) is the answer for f/g. If you wish you can factor the denominator: f/g = (x - 5)/{(x + 1)(x - 1) But either one should be ok. For the last one (e) just take the reciprocal of (d)

OpenStudy (anonymous):

can you show me the solution of (e)??

OpenStudy (ranga):

g/f = (x^2 - 1) / (x - 5) That is all.

OpenStudy (anonymous):

so the final answer for (d) is f/g = (x - 5)/{(x + 1)(x - 1)}

OpenStudy (ranga):

Yes. Either form is acceptable: f/g = (x - 5) / (x^2 - 1) or f/g = (x - 5) / { (x + 1)(x - 1) }

OpenStudy (anonymous):

ok , for (e) can you show me the solution??

OpenStudy (anonymous):

g/f=(x^2 - 1) / (x-5) ???

OpenStudy (ranga):

Yes. That is it.

OpenStudy (anonymous):

g/f={(x+1)(x-1)} / (x-5)

OpenStudy (ranga):

That is acceptable too. Either form is okay.

OpenStudy (anonymous):

thank you Ranga =)